OpenStudy (anonymous):

Which integer is closest to 0 on the number line? A. –19 B. 20 C. 32 D. 51

OpenStudy (swolesammy):

A

OpenStudy (rainbow_rocks03):

PLZ explain why it is a @swolesammy

OpenStudy (rainbow_rocks03):

what do you think it is @emiebug ?

OpenStudy (trojanpoem):

 To find the closest : ( number in the direction of inf - number in the other direction) 0 - (-19) = 19 units 20 - 0 = 20 units 32 - 0 = 32 units 51 - 0 = 51 units 19 is the least so A